OpenStudy (anonymous):

Find the unit rate if Mrs. Arnold drove 270 miles in 6 hours? Include details.

OpenStudy (ddcamp):

Unit rate is the amount of the one thing in 1 of the other things. An example (not this problem): 180 miles in 3 hours is 60 miles per hour

OpenStudy (anonymous):

so 270 divided by 6?

OpenStudy (texaschic101):

270/6 = x/1 (270 miles in 6 hours = x miles in 1 hour) cross multiply 270x = 6 x = 270/6 x = 45 Mrs. Arnold drove 45 miles in 1 hour (so unit rate is 45 mph )